# When Bias Pretends to Be Truth: How Spurious Correlations Undermine Hallucination Detection in LLMs

### v1

@@ Appendix C Neural Kernels and Kernel Regime
In the large width limit, the NTK converges to a deterministic kernel $\Theta$ and remains constant during training (yang2020neural; yang2021architectural). For the MSE loss $\ell(f(x),y)=\frac{1}{2}(f(x)-y)^{2}$, the solution of the kernel gradient descent has a closed form
@@
Both the NNGP and NTK kernels of neural networks are dot-product kernels, and indeed any dot-product kernel can be approximated by the NNGP or NTK of some neural network (simon2022reverse). Moreover, the neural kernels induced by general activations share the same properties as common kernels (e.g., Laplace kernels) through the norm equivalence of RKHS (holzmuller2025beyond).
@@ Appendix D Additional results
D.1 Hallucination Detection Algorithms
@@
Furthermore, we provide accuracy and TPR@5%FPR of detection methods for experiments in Section 3 (Figures 8 and 9). Across all evaluation metrics (accuracy, TPR@5%FPR, and AUROC in Section 3), performance consistently declines as $\rho$ increases, suggesting that spurious correlation systematically undermines hallucination detection methods.
@@
We only show the linear probing results for layer 21 as a representative example in the results above; detailed linear probing results of models trained from scratch and models finetuned from SmolLM2-1.7B under each $\rho$ are provided in Figures 10 and 11.
@@
D.2 Refusal Fine-tuning
@@
In Figures 13,14,15 we show the result that varies the correlation from $0.0$ to $0.9$ and each metric has the same meaning as before.

### v2

@@ Appendix C Neural Kernels
In the large width limit, the NTK converges to a deterministic kernel $\Theta$ and remains constant during training (lee2019wide; arora2019exact; yang2020neural; yang2021architectural). For the MSE loss $\ell(f(x),y)=\frac{1}{2}(f(x)-y)^{2}$, the solution of the kernel gradient descent has a closed form
@@
Equivalence to General Kernels.
Both the NNGP and NTK kernels of neural networks are dot-product kernels, and indeed any dot-product kernel can be achieved as the NNGP kernel or NTK of a suitably constructed neural network (simon2022reverse). Moreover, neural kernels derived from appropriately selected activation functions exhibit the same properties as a broad class of kernels through the norm equivalence of reproducing kernel Hilbert spaces (RKHS) (holzmuller2025beyond).
@@ Appendix D Additional results
D.1 Case Study of Hallucination in SimpleQA
In our analysis of the SimpleQA dataset, we observe numerous instances where hallucinations appear to be driven by spurious co-occurrence. Specifically, the model tends to output answers that have a strong statistical association (high Jaccard similarity) with entities in the question, even when those answers are factually incorrect.
To illustrate this phenomenon, we present a representative example involving an academic entity:
Question: To which academic society was computer scientist Sarita Vikram Adve elected in 2020?
Model Output: Association for Computing Machinery
Ground Truth: American Academy of Arts and Sciences
Although the model highly likely encountered the correct fact during pre-training [^1] https://en.wikipedia.org/wiki/Sarita_Adve, it fails to retrieve it. Instead, it outputs “Association for Computing Machinery” (ACM).
As analyzed in Table 2, this error aligns with the spurious correlation strength. The generic term “computer scientist” has a significantly higher Jaccard similarity with the hallucinated answer (0.0785) compared to the ground truth (0.0099). This suggests that the model falls back on the strong prior heuristic—Computer Scientists are often linked to ACM—overriding the specific factual constraint of the individual named in the prompt.
Table 2: Jaccard Similarity Analysis. We measure the co-occurrence strength between entities in the question and the answers. In this case, the hallucinated answer exhibits a much stronger correlation with the profession entity (“computer scientist”) than the ground truth does.
Question Entities
Answer Entities Generic: “computer scientist” Specific: “Sarita Vikram Adve”
Model Output (Hallucination):
Association for Computing Machinery 0.0785 0.0006
Ground Truth:
American Academy of Arts and Sciences 0.0099 0.0002
D.2 Spurious Correlation Induced by Style
In addition to the semantic spurious correlations discussed in the main text (i.e., synthetic surname-attribute mappings and real-world entity co-occurrence), we further investigate the impact of style correlation. Here, “style” refers to superficial textual properties—such as formality, emotional tone, or template structure—that are not causally related to the ground-truth answer but may act as heuristics for the model.
In real-world data, such correlations naturally arise; for instance, mathematical or scientific content is often concise and formal, whereas literary content tends to be narrative. If a model relies on these stylistic priors rather than factual reasoning, it may hallucinate when the prompt’s style superficially resembles contexts historically associated with a certain answer type.
Figure 7: Impact of Style Correlation on Detection Performance. We plot the AUROC of various hallucination detection methods across varying strengths of style correlation $\rho_{\text{style}}$.
Experimental Setup
To isolate this effect, we extend our synthetic data setting by introducing a controllable correlation between question templates and an auxiliary attribute (e.g., profession) during the Supervised Fine-Tuning (SFT) stage. Specifically, we control the correlation strength $\rho_{\text{style}}$: with probability $\rho_{\text{style}}$, a specific attribute (profession) is queried using a specific, fixed template structure; otherwise, the template is sampled uniformly from all available formats. For RFT, as in the previous method in section 3, we apply this template correlation only to the unknown person, while uniformly sampling templates for the known person. This creates a spurious correlation that the model might over- or under-reject, mistakenly relying on the question template. We maintain the same model architecture and training hyperparameters as in the main experiments.
Results of Detection Methods
As illustrated in Figure 7, while Perplexity initially achieves near-perfect performance (AUROC $\approx 1.0$) when correlation is absent, it suffers a severe collapse as the spurious style correlation strengthens; consistently, most other detection methods also exhibit a general performance degradation as the format correlation increases. This corresponds to the results in the Section 3.
Results of RFT
The experimental results are presented in Figure 8. We observe that stronger style correlations negatively impact model performance in two ways. First, for known individuals (whom the model should answer correctly), the QA accuracy declines as the correlation strength increases. Second, and more critically, the refusal mechanism for unknown individuals degrades: as the correlation intensifies, the model fails to explicitly reject unknown questions (i.e., the refusal rate drops), leading to increased hallucinations.
Figure 8: Impact of Style Correlation on RFT Performance. Left: Accuracy on Known Individuals: As the correlation strength increases, the model’s ability to correctly answer questions about known entities decreases. Right: Refusal Rate on Unknown Individuals: The model’s safety mechanism is compromised under strong correlation, resulting in a significant drop in refusal rate (i.e., the model fails to say “I don’t know” and instead hallucinates).
D.3 Hallucination Detection Algorithms
@@
Furthermore, we provide accuracy and TPR@5%FPR of detection methods for experiments in Section 3 (Figures 10 and 11). Across all evaluation metrics (accuracy, TPR@5%FPR, and AUROC in Section 3), performance consistently declines as $\rho$ increases, suggesting that spurious correlation systematically undermines hallucination detection methods.
@@
We only show the linear probing results for layer 21 as a representative example in the results above; detailed linear probing results of models trained from scratch and models finetuned from SmolLM2-1.7B under each $\rho$ are provided in Figures 12 and 13.
@@
D.4 Refusal Fine-tuning
@@
In Figures 15,16,17 we show the result that varies the correlation from $0.0$ to $0.9$ and each metric has the same meaning as before.